# LeoTCP: Low-Latency and High-Throughput Data Transport for LEO Satellite Networks

### v1

@@ Abstract
Low-Earth Orbit (LEO) satellite networks consist of thousands of satellites orbiting the Earth, enabling low-latency and high-throughput communications across the globe. Such networks present unprecedented challenges due to their dynamic nature, which state-of-the-art data transport protocols do not address. These include: (1) non-congestive latency variation and loss, caused by continuous satellite movement and fluctuating link quality due to weather effects; (2) transient hotspots leading to buffer build-up, latency inflation and, potentially, packet loss; and (3) frequent handovers, which may result in temporary connectivity loss and re-routing through paths with unknown congestion and delay characteristics. In this paper, we introduce LeoTCP, a novel data transport protocol designed specifically to target these challenges. LeoTCP leverages in-network telemetry (INT) to gather congestion information on a per-hop basis. Using this information, LeoTCP (1) minimises both buffer occupancy and latency for end-users (2) maximises application throughput and network utilisation, and (3) swiftly reacts to network hotspots. We compare LeoTCP to state-of-the-art data transport protocols using a LEO satellite simulation model and targeted micro-benchmarks, both based on OMNeT++/INET. The simulation model captures RTT dynamics in a simulated LEO satellite constellation, while the micro-benchmarks isolate key LEO specific characteristics, including non-congestive latency variation and loss, path changes, and congestion hotspots. Our results demonstrate that LeoTCP significantly increases goodput compared to existing state-of-the-art approaches while simultaneously minimising latency.
@@ I Introduction
The emergence of low Earth orbit (LEO) satellite networks, led by SpaceX’s Starlink [1] and followed by constellations from Amazon [2] and OneWeb [3], is transforming global connectivity. Each of these constellations will feature thousands of satellites that form a dynamic, interconnected mesh around the Earth. With an aggregate bandwidth projected to reach hundreds of terabits per second, they rival the total capacity of the global fiber infrastructure of today [4]. Inter-satellite links (ISLs) rely on free-space lasers, which transmit data at light speed in a vacuum, approximately 47% faster than through glass fiber [5]. LEO satellite networks promise high throughput and low latency, complementing the current terrestrial infrastructure for urban and, particularly, remote regions. However, to fully realise these benefits at the end-user level, network operators must carefully resource buffers and adopt transport protocols and congestion control (CC) mechanisms that can adapt to the highly dynamic characteristics of LEO networks and operate with low latency.
The dynamic nature of LEO satellite networks introduces new challenges that significantly affect the design of data transport protocols and CC mechanisms [6]. Satellites orbit at high speeds of 7.5km/s [7], and therefore ground stations (GSs) and user terminals (UTs) require frequent handovers, which may result in packet loss [8] and brief loss of connectivity. Such path changes can also lead to sudden and, potentially, substantial variations in the base RTT that network flows experience. Such RTT variations make it essential for transport protocols to maintain fairness between flows with different delays [9]. Additionally, shortest-path routing over the LEO mesh can lead to traffic hotspots [10], particularly along high-demand routes such as those between the US and Europe [11]. In addition to these issues, adverse weather conditions can degrade ground-to-satellite link quality, potentially introducing substantial non-congestive loss [12].
Traditional CC approaches, including specifically for satellite networks [13], do not consider such dynamic characteristics. Loss-based approaches operate with full queues [14, 6], undermining the inherent latency benefits of LEO satellite networks. Delay-based approaches frequently adjust transmission rates in response to satellite mobility-induced path reconfiguration [11, 6], resulting in buffer build-up or underutilisation [8]. NewReno and Cubic operate with substantially higher latencies [6], and perform poorly in environments with non-congestive loss. Other state-of-the-art approaches, such as BBR, aim to operate near the bandwidth-delay product (BDP) by estimating the bottleneck bandwidth and round-trip time (RTT) [15]. However, BBR relies on periodic probing to update its bandwidth and RTT estimates, which may fail to capture rapid path or latency changes, especially in LEO networks where such changes occur frequently, causing delayed adaptation until the next probe cycle. BBR has also been shown to be unfair when flows with different RTTs compete for the same bottleneck [16]. These limitations highlight the need for a data transport protocol that can react swiftly to LEO-specific dynamics and sustain high utilisation while keeping queue occupancy low. Recently, bespoke CC schemes tailored to LEO networks have begun to emerge. SaTCP [8] is an extension to TCP Cubic, which attempts to predict the occurrence of satellite handovers and route changes to minimise severe reactions to loss, which can occur due to handovers. However, it only attempts to mitigate the impact of these disruptive events; when no such event occurs, SaTCP behaves exactly like Cubic, with the issues outlined above. StarQUIC [17] is a similar approach that freezes the congestion window during handovers, for both BBRv3 and Cubic, and likewise shares the same issues.
((a)) Goodput
((b)) RTT
@@
To highlight the limitations of current data transport protocols under LEO dynamics, we simulate a single flow over a LEO satellite network path with periodic changes in bandwidth, RTT, and random non-congestive loss, updating every 15 seconds to match the rate at which Starlink UTs perform link reconfigurations [18]. [^1] The experimental setup and results are discussed further in Section IV-B. The solid black line in Figure 1(a) represents the optimal achievable throughput, while in Figure 1(b), it denotes the base RTT without any queuing delay. Figure 1(a) shows that Cubic achieves high throughput in the absence of loss (solid blue line) but induces significant latency inflation by persistently filling network buffers (shown in Figure 1(b)). Under non-congestive loss, Cubic’s performance collapses (dashed blue line). BBRv1, while more resilient to random loss (dashed green line), consistently falls short of fully utilising available bandwidth (solid green line). BBRv3 improves upon this (red solid line), showing better utilisation and somewhat reduced delay, but performs badly in the presence of non-congestive loss (dashed red line).
In this paper, we present LeoTCP, a transport protocol tailored for large-scale LEO satellite networks. LeoTCP is designed to operate under low buffer occupancy, enabling low-latency data transport while addressing key challenges unique to LEO environments that are inadequately handled by existing protocols. We employ in-network telemetry (INT) to obtain per-hop queue information, allowing the sender to adjust the congestion window according to the current network load. Our work builds on early work on INT-driven data transport in LEO satellite networks [9]. By utilising INT, we can address the main challenges associated with LEO satellite networks. (1) With INT information, we can determine the load at the bottleneck and adjust the sending rate to the calculated link capacity. We can minimise buffer occupancy, and, consequently, maintain low latency by adjusting LeoTCP’s sending rate to be just below the calculated bottleneck BDP. (2) LeoTCP can determine and quickly react to path changes, maximising goodput whilst minimising any potential loss. (3) INT information is used to quickly react to congestion in the network, reducing the potential for hotspots. (4) To minimise the negative effects of non-congestive loss [8, 12], and given that LeoTCP aims to operate with near-empty buffers at all times, packet loss is not used as an indication of congestion.
We implemented a simulation model of LeoTCP and compared its performance to state-of-the-art data transport protocols, namely TCP Cubic, BBR version 1, and BBR version 3. We used the OMNeT++ packet-level simulator [19] and the INET framework [20] to perform a series of microbenchmarks, comparing LeoTCP to the other approaches. These were designed to isolate and evaluate key characteristics of LEO satellite networks, including frequent path changes, variable RTTs, and the formation of congestion hotspots. These controlled tests provide insight into how such dynamics affect LeoTCP and state-of-the-art CC. We also employ a LEO satellite network model integrated with OMNeT++ and INET [21], which enables simulations with ISLs and bent-pipe (BP) links, allowing us to observe end-to-end latency variations and routing path shifts in a simulated LEO satellite constellation.
II Challenges and Objectives
@@
LEO satellite networks consist of large constellations of satellites orbiting at altitudes low enough to enable reduced latency compared to traditional geostationary systems. Their orbital parameters are defined by both altitude and inclination. For example, Starlink’s first orbital shell operates at an altitude of 550 km with a 53 inclination. Higher inclination angles, closer to 90, provide coverage near the poles, while lower inclinations primarily serve equatorial regions. At this altitude, satellites travel at $\sim$27,000km/h, resulting in short visibility windows and necessitating frequent handovers between satellites and ground interfaces [11, 8]. These networks rely on GSs, which serve as gateways to terrestrial networks, and UTs, which provide Internet access to end users. GSs are equipped with multiple phased array antennas, enabling simultaneous connections to several satellites [22]. In contrast, UTs are limited to a single-phased array antenna and can maintain only one active satellite link at a time. Early designs focused on BP routing, where ISLs were unavailable and satellites relayed data via GSs. BP paths can offer lower latency over short distances, especially when endpoints are near the same satellite [23]. However, the increasing adoption of inter-satellite links (ISLs) now enables traffic to be routed entirely through space. These ISLs support lower RTTs and higher throughput compared to traditional terrestrial paths [24].
@@
LEO satellite networks introduce a set of distinct challenges for data transport protocols and CC mechanisms. Unlike terrestrial networks, LEO constellations operate in a highly dynamic environment where the network topology and routing paths change constantly due to satellite mobility. This results in (1) non-congestive latency variation and loss, (2) transient hotspots and (3) frequent handovers.
@@
Non-congestive latency variation and loss. Propagation delays fluctuate due to satellite movement, leading to varying distances between UTs, GSs, and satellites. Furthermore, BP routing can amplify this variability, particularly over longer paths [23, 21]. This complexity is compounded by the structural mesh design of current LEO constellations, where shortest-path routing may not always yield optimal latency and often results in routing through congested or suboptimal links [25]. With high bandwidths and large variations of RTTs [11], buffer resourcing throughout the network must aim for minimal occupancy. In addition, rain fade can cause signal attenuation, resulting in slower data rates, increased latency, or temporary service interruptions [26]. These conditions can cause signal degradation, leading to higher latency, reduced throughput, and intermittent service interruptions [24].
@@
III LeoTCP Design
@@
LeoTCP employs INT and is built on top of TCP, allowing us to leverage TCP mechanisms such as Selective Acknowledgements (SACK) [28], timestamps [29], and the RACK loss detection algorithm [30]. The use of INT is inspired by existing congestion control approaches developed for high-throughput, low-latency environments [31], but involves major redesigns to ensure efficient and fair operation in LEO satellite networks. Switching devices operating on satellites and GSs generate INT information, which is appended to the header of each forwarded packet. Upon the packet’s arrival at the receiver, the telemetry data is extracted and embedded into the corresponding acknowledgement (ACK). With this information, LeoTCP senders can assess the bottleneck load on the current path and adapt their sending rates accordingly.
Below, we discuss the operational details of LeoTCP senders and switches, including the structure and usage of LeoTCP’s INT headers. LeoTCP receivers remain largely unmodified, with the only change being the relaying of received INT header data back to the sender into ACKs.
III-A LeoTCP Header
Figure 2 illustrates the LeoTCP INT header format, which is included in every packet. The fields pathID, baseRTT, and cwnd are inserted by the sender at transmission. These fields remain unchanged along the packet’s path through the network, except for the pathID. The pathID is updated at each hop and enables the sender to detect path changes (see Section III). cwnd is set by the sender to its current congestion window. baseRTT is set to the sender’s current estimate of the path’s base RTT. As in XCP [32], these values are used to compute the average RTT of all active flows traversing each switch. This calculation is stateless with respect to individual flows and is performed independently at each switch, enabling fair resource allocation, as discussed in Section III-C.
Figure 2: LeoTCP INT Header
TABLE I: LeoTCP INT Header Field Descriptions
@@ III-A LeoTCP Header
qLen Queue length at the time the packet leaves the buffer
rxQlen Queue length observed when the packet enters the buffer
@@
The additional fields shown in Figure 2 and described in Table I represent standard telemetry information that is appended by each switch along the path of the packet. The fields ts, qLen, txBytes, and bw provide information about the outgoing switch port. The ts field indicates the time at which the packet departed the queue, qLen represents the current buffer occupancy in bytes, txBytes denotes the total number of transmitted bytes, and bw is the bandwidth of the outgoing link. These fields are used to calculate the total utilisation of the outgoing link. Furthermore, avgRTT is computed using the baseRTT and cwnd fields (as mentioned above). Each switch sets the flowCnt value to its estimate of currently active LeoTCP flows; this can be done using the TCP four-tuple. To minimise network overhead, receivers could only echo back the fields needed by the LeoTCP sender to specify its sending rate based on the congestion at the current bottleneck; namely, ts, bw, qLen, rxQlen, txBytes, avgRTT, and flowCnt.
III-B LeoTCP Switch
Switches within the LEO satellite network append buffer load information to every outgoing packet via the INT header. Each switch maintains an average RTT (avgRTT), and the total number of active flows (flowCnt) for every interface. To track the number of flows, each switch identifies flows based on the TCP four-tuple. For the avgRTT value, each switch aggregates the baseRTTs of observed packets using a weighted average that accounts for their cwnd sizes. Specifically, it accumulates two values: the sum of baseRTTs weighted by the inverse of the cwnd, and the sum of squared baseRTTs also weighted by the inverse cwnd. Dividing these two values yields an approximation of the average RTT. These per-interval calculations can be trivially implemented in P4 programmable switches [33, 34]. Similarly to XCP [32], a timer set to the most recently observed average RTT triggers periodic updates of flowCnt and avgRTT. Switches append the total number of active flows (flowCnt) to the INT header, using the field of the same name, along with the per-hop congestion information. This enables LeoTCP senders to calculate their additive increase (AI) factor based on the level of congestion at the bottleneck, as described in Section III-C. Additionally, switches update the pathID in the INT header by XOR their own unique switchID with the current pathID. The final pathID is then used by the sender for detecting path changes, as described in Section III-C.
III-C LeoTCP Sender
@@
Per-RTT and Per-ACK Reactions. Our approach combines both per-RTT and per-ACK reactions, a strategy shown to enable rapid responses to congestion without triggering overreactions [31]. At the beginning of each sRTT, the sender sets the previous cwnd as a reference (refCwnd) for both additive and potential multiplicative decreases. The cwnd increases by one AI factor per sRTT. However, if overutilisation is detected, evaluated on a per-ACK basis, LeoTCP applies MD, scaled proportionally to the severity of congestion.
Target Utilisation. To estimate the level of bottleneck congestion, the sender computes a utilisation value, $U$, for each hop $i$ along the forward path, as in [31]. A $U$ value of 1 suggests that the link is at full capacity, while greater and lower than 1 implies over- and under-utilisation, respectively. The $U$ value is calculated as follows:
@@
This metric incorporates the INT information discussed in the previous section; specifically, it captures queue occupancy (qLen) and the outgoing rate ($txRate$) as a proportion of the link’s capacity (bw). The $txRate$ is defined as:
@@
To focus control on the most congested hop in its path, LeoTCP selects the one with the highest $U$ value as the current bottleneck and uses its INT information to adjust its window. To ensure stability and prevent overreaction to transient congestion, LeoTCP maintains an exponentially weighted moving average (EWMA) of the selected $U$ value.
@@
Dealing with latency variation. A key challenge in LEO satellite networks is the high variability in propagation delays. When computing the utilisation metric in Equation (1), using a sRTT measurement leads to unfairness across flows with varying latencies. Specifically, flows with higher RTTs will observe lower utilisation values for the same level of congestion, causing them to increase their sending rates more aggressively than lower RTT flows. CC algorithms that rely on feedback signals to manage transmission rates have been shown to become unstable with varying delays [37, 38]. In [31], the authors use a fixed base RTT when computing utilisation, assuming all flows experience the same delay, making it unsuitable for LEO satellite networks. To address this, LeoTCP uses an average RTT in Equation (1) which is computed by switches independently, as discussed in Section III-B. These values are appended to INT headers and echoed back to the sender, ensuring that all flows referencing the same bottleneck compute the same utilisation. In LeoTCP, the average RTT used in calculating utilisation ensures that window growth is scaled relatively to the bottleneck’s conditions, and independently of per-flow propagation delay, making it robust to the varying latencies found in LEO networks. [^2] In Section IV-E, we experimentally show the dramatic improvement in fairness that this design decision has in LeoTCP.
@@
A key limitation of using the instantaneous per-ACK RTT in INT headers is that they reflect both propagation and queueing delays. As queues build up along the path, the RTT increases, which in turn inflates the per-hop average RTT computed at switches. This has a compounding effect: the inflated RTT leads to a higher estimated BDP, prompting flows to set their cwnd higher than what is optimal, resulting in greater queue occupancy before overutilisation is detected. To mitigate this, LeoTCP leverages the per interface rxQlen and bw fields described in Section III-A, which represent the queue length before the packet is enqueued and the outgoing link bandwidth, respectively. By summing the ratio $\frac{\texttt{rxQlen}}{\texttt{bw}}$ across all hops, LeoTCP estimates the total queueing delay experienced along the path. Subtracting this from the instantaneous per-ACK RTT provides an approximation of the base RTT, being the propagation delay excluding queueing. To reduce noise and improve stability, the base RTT is smoothed using an EWMA. This base RTT is then appended by the sender to each INT header, allowing switches to compute an average base RTT across flows, rather than one influenced by queueing delays. By decoupling propagation delay from average RTT, this approach prevents cwnd from growing excessively in response to inflated per-ACK RTT values, helping to minimise persistent queue build-up and maintain low latency.
@@
Additive Increase and Multiplicative Decrease. LeoTCP employs an AIMD strategy based on the measured link utilisation U. AI is used to grow the cwnd; MD is applied only when utilisation exceeds the target $\eta=0.95$. This target is chosen to ensure high link utilisation while keeping buffer occupancy low. The AI factor is defined as follows, where bw and flowCnt is the bottlenecks values:
@@
AI is proportional to the estimated BDP of the link, which is computed using the sampled per-ACK RTT rather than the average. This ensures that flows with lower RTTs are not disadvantaged by slower growth. To ensure fair and BDP-proportional growth of flows, the bottleneck bandwidth extracted by the telemetry info is divided by the estimate of the number of active flows. This prevents queue build-up as the number of active flows changes. As discussed in Section III-B, switches maintain an estimate of the number of active flows, which senders use to compute the AI value.
When measured utilisation exceeds the target threshold $\eta$, LeoTCP applies an MD factor to reduce cwnd. The reduction factor is calculated as the ratio of the current utilisation U to the target $\eta$. This ensures that the reduction is proportional to the degree of overutilisation, meaning larger deviations from the target result in a more aggressive backoff. This design allows LeoTCP to maintain high link utilisation while responding swiftly to congestion and emerging hotspots. The AIMD scheme is described as follows, with the refCwnd being updated once every sRTT:
@@
Initial Phase. When new flows join a path already shared by many existing flows, they must be able to ramp up their sending rate quickly enough, whilst not saturating bottleneck buffers or hurting the performance of existing network flows. Since the AI factor is inversely proportional to the estimated number of flows, newly started flows may begin ramping up their sending rate too conservatively, delaying convergence to their fair share of bandwidth. To address this, LeoTCP introduces a bespoke start-up mechanism. When a new flow is created, it starts the initial phase. The flow then sends a small number of packets into the network to collect INT feedback. Once it receives the first INT header, the flow waits for an avgRTT (taken from the received INT header) before increasing its cwnd. This delay allows the switch-estimated flow count to reflect both the current flow, and, potentially, other newly arrived flows. After this waiting period, the sender computes its AI as 95% of the bottleneck BDP divided by the estimated flow count, rather than 5% in Algorithm 3. Using 95% as the AI allows newly arriving flows to converge rapidly to their fair rate while avoiding the need for unnecessary multiplicative decreases once the initial phase ends.

### v2

# OrbCC: High-Throughput and Low-Latency Data Transport for LEO Satellite Networks

@@ Abstract
The highly dynamic nature of Low-Earth Orbit (LEO) satellite networks introduces challenges that existing transport protocols fail to address, including non-congestive latency variation and loss, transient congestion hotspots, and frequent handovers that cause temporary disconnections and route changes with unknown congestion and delay characteristics. Our contention is that with this increase in complexity, there is insufficient information being returned from the network for existing congestion control algorithms to minimise latency while maintaining high throughput and minimising retransmissions. Our approach, OrbCC, leverages in-network support to collect per-hop congestion information and uses it to (1) minimise buffer occupancy and end-user latency, (2) maximise application throughput and network utilisation, and (3) rapidly respond to congestion hotspots. We evaluate OrbCC against state-of-the-art transport protocols using OMNeT++/INET-based LEO satellite simulations and targeted micro-benchmarks. The simulations capture RTT dynamics in a LEO constellation, while the micro-benchmarks isolate key characteristics such as non-congestive latency variation and loss, path changes, and congestion hotspots. Results show that OrbCC significantly improves goodput while simultaneously reducing latency and retransmissions compared to existing approaches.
@@ I Introduction
((a)) CDF of Goodput
((b)) CDF of RTT
((c)) CDF of Retransmissions
@@
Low Earth orbit (LEO) satellite networks consist of large constellations of satellites orbiting at altitudes low enough to enable reduced latency compared to traditional geostationary systems. Their orbital parameters are defined by both altitude and inclination. For example, Starlink’s first orbital shell operates at an altitude of 550 km with a 53 inclination. Higher inclination angles, closer to 90, provide coverage near the poles, while lower inclinations primarily serve equatorial regions. At this altitude, satellites travel at $\sim$7.5km/s, resulting in short visibility windows and necessitating frequent handovers between satellites and ground interfaces [4, 22, 5]. These networks rely on Ground Stations (GS), which serve as gateways to terrestrial networks, and User Terminals (UT) providing Internet access to end users. GSs are equipped with multiple phased array antennas, enabling simultaneous connections to several satellites [38]. In contrast, UTs are limited to a single-phased array antenna and can maintain only one active satellite link at a time. Early designs focused on Bent Pipe (BP) routing, where ISLs were unavailable and satellites relayed data via GSs. BP paths can offer lower latency over short distances, especially when endpoints are near the same satellite [16]. However, the increasing adoption of inter-satellite links (ISLs) now enables traffic to be routed entirely through space. These ISLs support lower RTTs and higher throughput compared to purely terrestrial paths [17].
@@
LEO satellite networks introduce a set of distinct challenges for data transport protocols and congestion control (CC) mechanisms. Unlike terrestrial networks, LEO constellations operate in a highly dynamic environment where the network topology and routing paths change constantly due to satellite mobility. This results in (1) non-congestive latency variation and loss, (2) transient hotspots and (3) frequent handovers.
@@
Non-congestive latency variation and loss. Propagation delays fluctuate due to satellite movement, leading to varying distances between UTs, GSs, and satellites. Furthermore, BP routing can amplify this variability, particularly over longer paths [16, 35]. This complexity is compounded by the structural mesh design of current LEO constellations, where shortest-path routing may not always yield optimal latency and often results in routing through congested or suboptimal links [2]. With high bandwidths and large variations of RTTs [22], buffer resourcing throughout the network must aim for minimal occupancy. In addition, rain fade can cause signal attenuation, resulting in slower data rates, increased latency, or temporary service interruptions [17, 27].
@@
Traditional CC approaches do not consider such dynamic characteristics. To highlight their limitations, we simulate a single flow over a LEO satellite network path with periodic changes in bandwidth, RTT, and random non-congestive loss, updating every 5 seconds to simulate rapid handovers (shown in Figure 1). [^1] The experimental setup and results are discussed further in Section III. Loss-based approaches such as Cubic operate with full queues [15, 1], undermining the inherent latency benefits of LEO satellite networks (shown in Figure 1(b)). Such approaches also cannot maintain high goodput due to the loss caused by handovers (solid blue line). Under non-congestive loss, Cubic’s performance collapses further (dashed blue line). Other state-of-the-art approaches, such as BBR, aim to operate near the bandwidth-delay product (BDP) by estimating the bottleneck bandwidth and round-trip time (RTT) [6]. However, BBR relies on periodic probing to update its bandwidth and RTT estimates, which may fail to capture rapid path or latency changes, especially in LEO networks where such changes occur frequently, causing delayed adaptation until the next probe cycle. BBR has also been shown to be unfair when flows with different RTTs compete for the same bottleneck [18]. BBRv1, while more resilient to random loss (dashed green line), consistently falls short of fully utilising available bandwidth (solid green line). BBRv3 slightly improves on this (solid red line), but likewise cannot update its model fast enough in the presence of rapid handovers. BBRv3 also suffers greatly with non-congestive loss (dashed red line). These limitations highlight the need for a data transport protocol that can react swiftly to LEO-specific dynamics and sustain high utilisation while keeping queue occupancy low. Recently, bespoke CC schemes tailored to LEO networks have begun to emerge. SaTCP [5] is an extension to Cubic, which attempts to predict the occurrence of satellite handovers and route changes to minimise severe reactions to loss. However, it only attempts to mitigate the impact of these disruptive events; when no such event occurs, SaTCP behaves exactly like Cubic, with the issues outlined above (dashed purple line). StarQUIC [21] is a similar approach that freezes the congestion window during handovers, for both BBRv3 and Cubic, and likewise shares the same issues. LeoCC [25] is another recent CC algorithm, which extends BBRv1 and attempts to detect path reconfigurations, allowing it to refresh its bottleneck model to avoid relying on outdated bandwidth/RTT samples. While effective in the scenarios it targets, LeoCC incurs repeated startup phases for every reconfiguration, leading to spikes of congestion and loss after each reconfiguration (see Figure 1(c)). LeoCC additionally relies on ICMP probing every 10ms to detect reconfigurations, with the assumption that there is a separate queue management for ICMP and TCP/UDP traffic. This assumption is validated for Starlink but may not hold for future or non-Starlink deployments.
Our contention is that by increasing the information about network state being returned to senders through in-network telemetry (INT), CC algorithms can ensure low buffer occupancy and high goodput while avoiding unnecessary retransmissions, thereby preserving the performance potential of LEO networks. Motivated by this, we propose OrbCC, which leverages per-hop, switch-reported signals to obtain queue information, allowing the sender to adjust the congestion window according to the current network load. Naively applying existing terrestrial CC designs that employ in-network support [26] is insufficient in LEO satellite networks due to the LEO environment violating their key assumptions. OrbCC instead directly tackles the LEO-specific challenges: (1) We can determine the load at the bottleneck and adjust the sending rate to the calculated link capacity. We can minimise buffer occupancy, and, consequently, maintain low latency by adjusting OrbCC’s sending rate to be just below the calculated bottleneck BDP. (2) OrbCC can determine and quickly react to path changes, maximising goodput whilst minimising any potential loss. (3) We can quickly react to congestion in the network, reducing the potential for hotspots. (4) To minimise the negative effects of non-congestive loss [5, 36], and given that OrbCC aims to operate with near-empty buffers at all times, packet loss is not used as an indication of congestion. Building on this, OrbCC introduces three novel mechanisms validated in our evaluation: (i) rapid path-change detection and reaction to sustain goodput across handovers, (ii) improved fairness across diverse path delays, and (iii) an initial-phase strategy that enables new flows to ramp up rapidly and fairly.
II OrbCC Design
@@
OrbCC is built on top of TCP, allowing us to leverage TCP mechanisms such as Selective Acknowledgements (SACK) [10], timestamps [20], and the RACK loss detection algorithm [7]. Switching devices, operating on satellites and GSs, report per-hop network state information, which is appended to the header of each forwarded packet. The use of such in-network support has been explored in the past in the context of data centers [26], but involves major redesigns to ensure efficient and fair operation in LEO satellite networks. Upon the packet’s arrival at the receiver, the hop data is extracted and embedded into the corresponding acknowledgement (ACK). With this information, OrbCC senders can assess the bottleneck load on the current path and adapt their sending rates accordingly.
Each switch maintains an average RTT (avgRTT), and the total number of active flows (flowCnt) for every interface. To track the number of flows, each switch identifies flows based on the TCP four-tuple. For the avgRTT value, each switch aggregates the baseRTTs of observed packets using a weighted average that accounts for their cwnd sizes. Specifically, it accumulates two values: the sum of baseRTTs weighted by the inverse of the cwnd, and the sum of squared baseRTTs also weighted by the inverse cwnd. Dividing these two values yields an approximation of the average RTT. These per-interval calculations can be trivially implemented in P4 programmable switches [3, 24]. Similarly to XCP [23], a timer set to the most recently observed average RTT triggers periodic updates of flowCnt and avgRTT. Switches append flowCnt to the header, using the field of the same name, along with the per-hop congestion information. This enables OrbCC senders to calculate their additive increase (AI) factor based on the level of congestion at the bottleneck, as described below. Additionally, switches update the pathID in the header by XOR-ing their own unique switchID with the current pathID. The final pathID is then used by the sender for detecting path changes.
Figure 2 illustrates the OrbCC header format, which is included in every packet. At transmission, the sender inserts pathID, baseRTT, and cwnd, resulting in approximately 52 bits of fixed per-packet overhead. These fields remain unchanged along the path, except as described above for pathID.
Figure 2: OrbCC Header
TABLE I: OrbCC Header Field Descriptions
@@
qLen Queue length (bytes) when the packet leaves the buffer
rxQlen Queue length (bytes) when the packet enters the buffer
@@
The additional fields shown in Figure 2 and described in Table I represent the information appended by each switch along the path of the packet. The fields ts, qLen, rxQlen, txBytes, and bw provide information about the outgoing switch port and are used to calculate the total utilisation of the outgoing link. Each switch sets the flowCnt value to its estimate of currently active OrbCC flows using the TCP four-tuple. These per-hop fields contribute approximately 132 bits ($\approx$17 bytes) of overhead. An average hop count among major city pairs is approximately 10 [2], leading to roughly 170 bytes of overhead per packet. This represents an average upper bound, since the packet is largest only near the end of the path. In practice, this overhead is modest relative to MSS-sized packets and is outweighed by OrbCC’s goodput gains from precise congestion feedback. To minimise network overhead, receivers only echo back the fields corresponding to the identified current bottleneck switch.
@@
Per-RTT and Per-ACK Reactions. Our approach combines both per-RTT and per-ACK reactions, a strategy shown to enable rapid responses to congestion without triggering overreactions. At the beginning of each sRTT window, the sender sets the previous cwnd as a reference (refCwnd) for both additive and potential multiplicative decreases. The cwnd increases by one AI factor per sRTT. However, if overutilisation is detected, evaluated on a per-ACK basis, OrbCC applies MD, scaled proportionally to the severity of congestion.
Target Utilisation. To estimate the level of bottleneck congestion, the sender computes a utilisation value, $U$, for each hop $i$ along the forward path. A $U$ value of 1 suggests that the link is at full capacity, while greater and lower than 1 implies over- and under-utilisation, respectively. The $U$ value is calculated as follows:
@@
This metric uses the hop information to capture the queue occupancy (qLen) and the outgoing rate ($txRate$) as a proportion of the link’s capacity (bw). The $txRate$ is defined as:
@@
To focus control on the most congested hop in its path, OrbCC selects the one with the highest $U$ value as the current bottleneck and uses the respective network provided information to adjust its window. To ensure stability and prevent overreaction to transient congestion, OrbCC maintains an exponentially weighted moving average (EWMA) of the selected $U$ value.
@@
Dealing with latency variation. A key challenge in LEO satellite networks is the high variability in propagation delays. When computing the utilisation metric in Equation (1), using a sRTT measurement leads to unfairness across flows with varying latencies. Specifically, flows with higher RTTs will observe lower utilisation values for the same level of congestion, causing them to increase their sending rates more aggressively than lower RTT flows. CC algorithms that rely on feedback signals to manage transmission rates have been shown to become unstable with varying delays [28, 33]. To address this, OrbCC uses an average RTT in Equation (1) which is computed by switches independently, as described above. These values are appended to the headers and echoed back to the sender, ensuring that all flows referencing the same bottleneck compute the same utilisation. In OrbCC, the average RTT used in calculating utilisation ensures that window growth is scaled relative to the bottleneck’s conditions, and independently of per-flow propagation delay, making it robust to the varying latencies found in LEO networks. [^2] In Section III, we experimentally show the dramatic improvement in fairness that this design decision has in OrbCC.
@@
A key limitation of using the instantaneous per-ACK RTT measurements is that they reflect both propagation and queueing delays. As queues build up along the path, the RTT increases, which in turn inflates the per-hop average RTT computed at switches. This has a compounding effect: the inflated RTT leads to a higher estimated BDP, prompting flows to set their cwnd higher than what is optimal, resulting in greater queue occupancy before overutilisation is detected. To mitigate this, OrbCC leverages the per-interface rxQlen and bw fields, which represent the queue length before the packet is enqueued and the outgoing link bandwidth, respectively. By summing the ratio $\frac{\texttt{rxQlen}}{\texttt{bw}}$ across all hops, OrbCC estimates the total queueing delay experienced along the path. Subtracting this from the instantaneous per-ACK RTT provides an approximation of the base RTT. To reduce noise, the base RTT is smoothed using an EWMA and appended to each header, enabling switches to compute an average base RTT across flows independent of queueing delays. This decoupling of propagation delay from RTT prevents excessive cwnd growth due to inflated per-ACK RTTs, helping to minimise persistent queue build-up and maintain low latency.
@@
Additive Increase and Multiplicative Decrease. OrbCC employs an AIMD strategy based on the measured link utilisation U. AI is used to grow the cwnd; MD is applied only when utilisation exceeds the target $\eta=0.95$. This target is chosen to ensure high link utilisation while keeping buffer occupancy low. The AI factor is defined as follows, where bw and flowCnt are the bottleneck’s values:
@@
AI is proportional to the estimated BDP of the link, which is computed using the sampled per-ACK RTT rather than the average. This ensures that flows with lower RTTs are not disadvantaged by slower growth. To ensure fair and BDP-proportional growth of flows, the bottleneck bandwidth is divided by the estimate of the number of active flows. This prevents queue build-up as the number of active flows changes. As described above, switches maintain an estimate of the number of active flows, which senders use to compute the AI value.
When measured utilisation exceeds the target threshold $\eta$, OrbCC applies an MD factor to reduce cwnd. The reduction factor is calculated as the ratio of the current utilisation U to the target $\eta$. This ensures that the reduction is proportional to the degree of overutilisation, meaning larger deviations from the target result in a more aggressive backoff. This design allows OrbCC to maintain high link utilisation while responding swiftly to congestion and emerging hotspots. The AIMD scheme is described as follows, with the refCwnd being updated once every sRTT:
@@
Initial Phase. When new flows join a path already shared by many existing flows, they must be able to ramp up their sending rate quickly enough, whilst not saturating bottleneck buffers or hurting the performance of existing network flows. Since the AI factor is inversely proportional to the estimated number of flows, newly started flows may begin ramping up their sending rate too conservatively, delaying convergence to their fair share of bandwidth. To address this, OrbCC introduces a bespoke start-up mechanism. When a new flow is created, it starts the initial phase. The flow then sends a small number of packets into the network to collect hop feedback. Once it receives the first header, the flow waits for an avgRTT (taken from the respective header) before increasing its cwnd. This delay allows the switch-estimated flow count to reflect both the current flow, and, potentially, other newly arrived flows. After this waiting period, the sender computes its AI as 95% of the bottleneck BDP divided by the estimated flow count, rather than 5% in Algorithm 3. Using 95% as the AI allows newly arriving flows to converge rapidly to their fair rate while avoiding the need for unnecessary multiplicative decreases once the initial phase ends.